OpenStudy (mybarra):

pi/3simpi/6=1/2(simpi/2-simpi/6)

OpenStudy (mathmale):

sorry to have to ask, but could you please draw this problem in the Draw utility (below)? In my 72 years I've never seen an expression such as pi/3simpi/6. Please explain. Better yet, present this problem as a drawing.

jimthompson5910 (jim_thompson5910):

pi/3simpi/6=1/2(simpi/2-simpi/6) pi/3 sim pi/6=1/2( sim pi/2-sim pi/6) pi/3 \sim pi/6=1/2( \sim pi/2-\sim pi/6) \[\Large pi/3 \sim pi/6=1/2( \sim pi/2-\sim pi/6)\] That's just a guess at what @mybarra meant to write. Though I'm not sure why you'd want to use the \(\Large \sim\) operator

jimthompson5910 (jim_thompson5910):

or did you mean to say \(\Large \text{sin}\) instead of \(\Large \text{sim}\) ?

OpenStudy (mathmale):

@mybarra: Please, would you clear up what you meant? Often others on OpenStudy can figure out what is being asked, but why impose additional work on potential helpers?